OpenStudy (5335202):

L is the midpoint of CD. If CL= 1/3x+8 and LD=2/3x-4, find the length of CD. Would i add the equations together?? HELP

OpenStudy (anonymous):

L divides CD in half, meaning CL and LD would be exactly the same.

OpenStudy (anonymous):

So CL=LD \[\frac{ 1 }{ 3x+8 }=\frac{ 2 }{ 3x-4 }\]

OpenStudy (anonymous):

Then you would cross multiply. 1(3x-4) = 2(3x+8) Tell me what you get ^^

OpenStudy (5335202):

3x-4=6x+16??

OpenStudy (anonymous):

Yep ^^ Now combine like terms

OpenStudy (5335202):

ooohhhhhh -4=3x+16 -20=3x -6.66=x right?

OpenStudy (anonymous):

Yep, but it'll be easier to work if you keep it as a fraction. so now that you know x is -20/3, plug -20/3 into either CL or LD to find its length

OpenStudy (5335202):

ooohhhh thank you so much!

OpenStudy (anonymous):

Mhm. And once you get the length, remember to multiply it by 2 because you would only be finding half the length by using CL or LD instead of CD

OpenStudy (5335202):

Thank you! i never would've done that!